# supplement

### v1

@@ 2.1 A review of ABC-MCMC
where $\boldsymbol{x}$ denotes synthetic data simulated from $p(\cdot\mid{\boldsymbol{\theta}})$, $K_{\varepsilon}(\cdot,\cdot)$ is a kernel function and the threshold $\varepsilon$ plays a role of bandwidth. The marginal posterior distribution $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})$ can be evaluated by integrating out the auxiliary data set $\boldsymbol{x}$ from $\pi_{\varepsilon}({\boldsymbol{\theta}},\boldsymbol{x}\mid\boldsymbol{y})$. The choice of threshold $\varepsilon$ is a trade-off between accuracy and computational speed. A smaller value of threshold can lead to more accurate ABC posterior, but with higher computational cost. When the threshold value $\varepsilon$ approaches to zero, $\pi_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})$ converges to the exact Bayesian posterior $\pi(\boldsymbol{y}\mid{\boldsymbol{\theta}})$.
To sample from high-dimensional posterior space, rejection ABC is not efficient since there is often a significant mismatch between the prior distribution and the posterior. In some cases, generating millions of synthetic data sets may be required to obtain a few accepted samples. The ABC-MCMC approach is a good alternative to rejection ABC for likelihood-free inference. It uses local moves for proposing parameters, and is thus more efficient. The idea for MCMC is to construct a Markov chain whose probability density converges to the stationary distribution $\pi_{\varepsilon}({\boldsymbol{\theta}},\boldsymbol{x}\mid\boldsymbol{y})$, with the marginal distribution being $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})$. We introduce a proposal distribution $q({\boldsymbol{\theta}}^{*}\mid{\boldsymbol{\theta}})$ and a function $\alpha$ that satisfies
$\pi_{\varepsilon}({\boldsymbol{\theta}},\boldsymbol{x}\mid\boldsymbol{y})q({\boldsymbol{\theta}}^{*}\mid{\boldsymbol{\theta}})p(\boldsymbol{x}^{*}\mid{\boldsymbol{\theta}}^{*})\alpha\left[({\boldsymbol{\theta}},\boldsymbol{x}),({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*})\right]=\pi_{\varepsilon}({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*}\mid\boldsymbol{y})q({\boldsymbol{\theta}}\mid{\boldsymbol{\theta}}^{*})p(\boldsymbol{x}\mid{\boldsymbol{\theta}})\alpha\left[({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*}),({\boldsymbol{\theta}},\boldsymbol{x})\right],$
where
$\alpha\left[({\boldsymbol{\theta}},\boldsymbol{x}),({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*})\right]=\operatorname{min}\left\{1,\frac{\pi({\boldsymbol{\theta}}^{*})K_{\varepsilon}(\boldsymbol{x}^{*},\boldsymbol{y})q({\boldsymbol{\theta}}\mid{\boldsymbol{\theta}}^{*})}{\pi({\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})q({\boldsymbol{\theta}}^{*}\mid{\boldsymbol{\theta}})}\right\}.$ (2.1)
At $(t+1)$-th MCMC iteration, we repeat the following procedure.
  1. (i)
    Sample ${\boldsymbol{\theta}}^{*}$ from $q(\cdot\mid{\boldsymbol{\theta}}_{t})$.
  2. (ii)
    Generate synthetic data $\boldsymbol{x}^{*}$ from $P_{{\boldsymbol{\theta}}^{*}}$.
  3. (iii)
    Accept $({\boldsymbol{\theta}}_{t+1},\boldsymbol{x}_{t+1})=({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*})$ with probability $\alpha\left[({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t}),({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*})\right]$; otherwise, set $({\boldsymbol{\theta}}_{t+1},\boldsymbol{x}_{t+1})=({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$.
2.2 An ABC-MCMC with global and local proposals
The exploration of the standard ABC-MCMC introduced in Section 2.1 is constrained to a small region around the current state. This often results in poor mixing, as particles can become trapped in some local regions, especially for multi-modal ABC posteriors. Therefore, it is necessary to design more global updates to improve the exploration ability of MCMC, helping samples move across different modes more frequently. However, globally proposing each parameter may not adequately capture the dependencies and correlations between parameters, making it challenging to explore the space effectively. By integrating both local and global proposals, the algorithm can combine the strengths of both approaches, achieving an optimal balance that allows for a detailed examination of local regions while ensuring a more comprehensive exploration of the entire parameter space.
We implement the Global-Local ABC-MCMC algorithm by alternating between local and global sampling. Specifically, we perform global sampling with probability $\gamma$ and local sampling with probability $1-\gamma$ at each iteration. The transition kernel of the Global-Local ABC-MCMC is defined as:
$k((\boldsymbol{\theta},\mathbf{x}),(\boldsymbol{\theta}^{*},\mathbf{x}^{*}))=\gamma\,k_{g}((\boldsymbol{\theta},\mathbf{x}),(\boldsymbol{\theta}^{*},\mathbf{x}^{*}))+(1-\gamma)\,k_{l}((\boldsymbol{\theta},\mathbf{x}),(\boldsymbol{\theta}^{*},\mathbf{x}^{*})),$ (2.2)
where $k_{l}$ is a local ABC-MCMC kernel, and $k_{g}$ is a global ABC-MCMC kernel. This formulation provides flexibility in the explore-exploit strategy by incorporating both local and global perspectives. The parameter $\gamma$ controls the trade-off between exploitation and exploration. When $\gamma=0$, the Global-Local ABC-MCMC algorithm degenerates to the standard ABC-MCMC; otherwise, when $\gamma=1$, it only utilizes a global proposal.
2.3 The global proposal: iteration sampling importance resample
One straightforward approach is to use a proposal distribution $q(\cdot)$ that is independent of the previous state, serving as a global proposal. This can facilitate more frequent transitions between modes. However, with such an independent proposal, the MH acceptance probability $\alpha\left[({\boldsymbol{\theta}},\boldsymbol{x}),({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*})\right]=\operatorname{min}\left\{1,\frac{\pi({\boldsymbol{\theta}}^{*})K_{\varepsilon}(\boldsymbol{x}^{*},\boldsymbol{y})q({\boldsymbol{\theta}})}{\pi({\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})q({\boldsymbol{\theta}}^{*})}\right\}$ tends to be very low if $q(\cdot)$ diverges significantly from the ABC posterior. This presents challenges for the global exploration capabilities of GL-ABC-MCMC, particularly in high-dimensional spaces. Therefore, it is essential for the global proposal to be as close to the ABC posterior as possible. However, inferring this proposal is nearly as complex as inferring the ABC posterior itself. In response to this challenge, we propose a flexible approach to obtain an efficient global proposal.
Sampling Importance Resampling (SIR) is a commonly used technique to transition from a simple distribution to a more complex one. It involves obtaining samples from a target distribution by resampling from a set of samples with weights. These samples are drawn from a proposal distribution, and the weights are adjusted based on the ratio of the target distribution to the proposal distribution. Thus, we could utilize iterative SIR (i-SIR) to design efficient global proposal distributions. The likelihood-based i-SIR was originally proposed in Tjelmeland (2004) and further studied in Andrieu et al. (2010, 2018); Samsonov et al. (2022). Here, we introduce a likelihood-free i-SIR as a more efficient global sampler for ABC-MCMC, referred to as ABC-i-SIR. Algorithm 1 illustrates a single stage of ABC-i-SIR. In one stage of ABC-i-SIR, we sample $\left\{{\boldsymbol{\theta}}_{(j)}^{*}\right\}_{j=1:N_{b}}$ from the independent proposal distribution, then generate corresponding synthetic data $\left\{\boldsymbol{x}_{(j)}^{*}\right\}_{j=1:N_{b}}$. We then combine the newly generated particles with the last state of the chain $({\boldsymbol{\theta}}_{(0)}^{*},\boldsymbol{x}_{(0)}^{*})=({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$ as candidate particles. The new state $({\boldsymbol{\theta}}_{t+1},\boldsymbol{x}_{t+1})$ is obtained by sampling from the candidate particles $\left\{{\boldsymbol{\theta}}_{(j)}^{*},\boldsymbol{x}_{(j)}^{*}\right\}_{j=0:N_{b}}$ with importance weights $w_{j}\propto\pi({\boldsymbol{\theta}}_{(j)}^{*})K_{\varepsilon}(\boldsymbol{x}_{(j)}^{*},\boldsymbol{y})/q({\boldsymbol{\theta}}_{(j)}^{*})$. In Proposition 1 and 2 of Section 2.5, we show that the ABC-i-SIR algorithm can be interpreted as a systematic-scan two-stage Gibbs sampler and demonstrate the uniform geometric ergodicity of the ABC-i-SIR kernel.
Algorithm 1 Single stage of ABC-i-SIR algorithm
Procedure ABC-i-SIR($({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$, $q$, $\pi$, $P_{\boldsymbol{\theta}}$, $K_{\varepsilon}$)
@@ 2.3 The global proposal: iteration sampling importance resample
Output: New state $({\boldsymbol{\theta}}_{t+1}$,$\boldsymbol{x}_{{t+1}})$.
@@
2.4 Optimization of hyper-parameters

### v2

@@ 2.2 The global move based on iterated-sampling importance resample
Output: New state $({\boldsymbol{\theta}}_{t+1}$$\boldsymbol{x}_{{t+1}})$.
@@
A simple choice of the importance proposal in ABC-i-SIR is the prior distribution. For complex models, to improve algorithmic efficiency, we can select a larger threshold $\varepsilon^{*}>\varepsilon$ and perform a small number of simulations to obtain a rough estimate of the posterior $\pi_{\varepsilon^{*}}$, which is then used as the importance proposal for ABC-i-SIR.
In Proposition 1 of Section 4, we show that the ABC-i-SIR algorithm can be interpreted as a systematic-scan two-stage Gibbs sampler. Given that candidate parameters are drawn from a proposal distribution independent of the current state, each iteration of ABC-i-SIR effectively performs a global ABC-MCMC update. Moreover, Proposition 2 in Section 4 establishes the uniform geometric ergodicity of the MCMC kernel induced by ABC-i-SIR.
2.3 A gradient based local proposal
In practice, it is common to employ a $p$-dimensional normal distribution as proposal for local MCMC steps (e.g. $q({\boldsymbol{\theta}}^{*}\mid{\boldsymbol{\theta}})=\mathcal{N}({\boldsymbol{\theta}}^{*};{\boldsymbol{\theta}},\sigma^{2}\boldsymbol{I}_{p})$). The scale parameter of the normal distribution significantly influences the efficiency and performance of the algorithm, particularly for high-dimensional parameter space. Designing an appropriate scale poses a challenge.
It’s necessary to design a general proposal mechanism which provides large proposal transitions with high probability of acceptance. Metropolis adjusted Langevin algorithm (MALA) (Grenander1994MALA; roberts1996exponential; roberts1998optimal; girolami2011riemann) combines Langevin dynamics with a Metropolis-Hastings correction to create an adaptive and efficient MCMC algorithm.
The Langevin diffusion process is defined by the stochastic differential equation (SDE)
$d{\boldsymbol{\theta}}(t)=\nabla_{\boldsymbol{\theta}}\mathcal{L}\{{\boldsymbol{\theta}}(t)\}dt/2+d\boldsymbol{W}(t),$ (1)
where $\mathcal{L}({\boldsymbol{\theta}})$ denotes the logarithm of target distribution, $\boldsymbol{W}$ denotes a $p$-dimensional Brownian motion. The diffusion is irreducible, strong Feller and aperiodic, with stationary distribution $\pi_{T}(\cdot)\propto\exp(\mathcal{L}(\cdot))$. This process can be regarded as a continuous-time sampling method. Unfortunately, the implementation is infeasible in practice. The common method is to use discretized approximations via Euler-Maruyama discretization:
${\boldsymbol{\theta}}^{*}={\boldsymbol{\theta}}+\eta^{2}\nabla_{\boldsymbol{\theta}}\mathcal{L}({\boldsymbol{\theta}})/2+\eta z,$ (2)
where $\eta$ is the step-size and $z\sim\mathcal{N}(0,\boldsymbol{I}_{p})$. However, when discretizing the diffusion, some bias is introduced, then convergence to the invariant distribution $\pi_{T}$ is no longer guaranteed. MALA employs a Metropolis-Hastings rejection-acceptance step after every iteration to ensure the convergence property. The discrete form of Langevin diffusion plays the role of an instrumental proposal distribution. In particular, $q\left(\boldsymbol{\theta}^{*}\mid\boldsymbol{\theta}\right)=\mathcal{N}\left\{\boldsymbol{\theta}\mid{\boldsymbol{\theta}}+\eta^{2}\nabla_{\boldsymbol{\theta}}\mathcal{L}({\boldsymbol{\theta}})/2,\eta^{2}\boldsymbol{I}_{p}\right\}$ and the corresponding acceptance probability is $\alpha({\boldsymbol{\theta}},{\boldsymbol{\theta}}^{*})=\min\left\{1,\frac{\exp(\mathcal{L}({\boldsymbol{\theta}}^{*}))p({\boldsymbol{\theta}}\mid{\boldsymbol{\theta}}^{*})}{\exp(\mathcal{L}({\boldsymbol{\theta}}))p({\boldsymbol{\theta}}^{*}\mid{\boldsymbol{\theta}})}\right\}$.
Recall that our target distribution is $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})\propto\pi({\boldsymbol{\theta}})p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})$, where $p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})=\int p(\boldsymbol{x}\mid{\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})d\boldsymbol{x}$ is the approximate likelihood. Then $\mathcal{L}({\boldsymbol{\theta}})=\log(\pi({\boldsymbol{\theta}}))+\log(p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}}))$. The core of ABC-MALA is computing the gradient of log-likelihood. meeds2015HABC apply HMC to ABC and propose to compute the gradient using common random numbers (CRN). The key idea of CRN is to represent $\boldsymbol{x}$ as a deterministic function of variables $\boldsymbol{\theta}$} and $\omega$, that is to define the random variable $\boldsymbol{x}$ from $p(\boldsymbol{x}\mid{\boldsymbol{\theta}})$ as a deterministic function $f({\boldsymbol{\theta}},\omega)$, where $\omega$ is a random seed. Then, the estimated likelihood based on $\varepsilon$-kernel is
$p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})=\sum_{s=1}^{S}K_{\varepsilon}(f({{\boldsymbol{\theta}}},\omega_{s}),\boldsymbol{y}),$ (3)
where $\omega_{s}$ ($s=1,\ldots,S$) are $S$ random seeds. meeds2015HABC suggest to approximate the conditional distribution $p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})$ by $K_{\varepsilon}(f({\boldsymbol{\theta}},\omega_{m}),\boldsymbol{y})$ under the assumption that the threshold is very close to 0. Here, $\omega_{m}$ represents a seed that ensures $K_{\varepsilon}(f({\boldsymbol{\theta}},\omega_{m}),\boldsymbol{y})$ is the largest among $\left\{K_{\varepsilon}(f({\boldsymbol{\theta}},\omega_{i}),\boldsymbol{y})\right\}_{i=1:S}$. Then, the gradient of log-likelihood is estimated by
$\nabla_{\boldsymbol{\theta}}\log(p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}}))=\frac{1}{2d_{\boldsymbol{\theta}}}\left\{\log(K_{\varepsilon}(f({{\boldsymbol{\theta}}}+d_{\boldsymbol{\theta}},\omega_{m}^{+}),\boldsymbol{y}))-\log(K_{\varepsilon}(f({{\boldsymbol{\theta}}}-d_{\boldsymbol{\theta}},\omega_{m}^{-}),\boldsymbol{y}))\right\},$ (4)
where $\omega_{m}^{+}=\arg\max_{\omega_{i}\in\{\omega_{1:S}\}}K_{\varepsilon}(f({{\boldsymbol{\theta}}}+d_{\boldsymbol{\theta}},\omega_{i}),\boldsymbol{y})$ and $\omega_{m}^{-}=\arg\max_{\omega_{i}\in\{\omega_{1:S}\}}K_{\varepsilon}(f({{\boldsymbol{\theta}}}-d_{\boldsymbol{\theta}},\omega_{i}),\boldsymbol{y})$. In this article, we denote the method as $\text{CRN}_{max}$. In practice, ensuring the efficiency of ABC methods with a very small $\varepsilon$ is challenging. Here, we estimate the partial derivative with respect to $\boldsymbol{\theta}$} without the small threshold assumption as follows:
$\nabla_{\boldsymbol{\theta}}\log(p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}}))=\frac{1}{2d_{\boldsymbol{\theta}}}\left(\log\left(\sum_{s=1}^{S}K_{\varepsilon}(f({{\boldsymbol{\theta}}}+d_{\boldsymbol{\theta}},\omega_{s}),\boldsymbol{y})\right)-\log\left(\sum_{s=1}^{S}K_{\varepsilon}(f({{\boldsymbol{\theta}}}-d_{\boldsymbol{\theta}},\omega_{s}),\boldsymbol{y})\right)\right),$ (5)
where $d_{\boldsymbol{\theta}}$ is a small perturbation. The detailed finite difference stochastic approximation with CRN ($\text{CRN}_{mean}$) is shown in the Appendix A.1.
However, in the tail region of the posterior distribution, the logarithmic function can exacerbate the errors associated with Monte Carlo integration, leading to significantly inaccurate gradient estimates. In this study, we employ a Gaussian distribution to model the likelihood based on simulated data to obtain a more stable and accurate estimator when there are few simulations to estimate the integral. By utilizing the Gaussian kernel as the kernel density function, we derive the gradient estimate as follows:
$\nabla_{\boldsymbol{\theta}}\log(p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}}))=\frac{1}{2d_{\boldsymbol{\theta}}}\left(-\frac{1}{2}\log\left(\frac{\hat{\sigma}_{{\boldsymbol{\theta}}^{+}}^{2}+\varepsilon^{2}}{\hat{\sigma}_{{\boldsymbol{\theta}}^{-}}^{2}+\varepsilon^{2}}\right)-\frac{(\boldsymbol{y}-\hat{\mu}_{{\boldsymbol{\theta}}^{+}})^{2}}{2(\hat{\sigma}_{{\boldsymbol{\theta}}^{+}}^{2}+\varepsilon^{2})}+\frac{(\boldsymbol{y}-\hat{\mu}_{{\boldsymbol{\theta}}^{-}})^{2}}{2(\hat{\sigma}_{{\boldsymbol{\theta}}^{-}}^{2}+\varepsilon^{2})}\right),$ (6)
3 Two adaptive schemes
3.1 An adaptive proposal of ABC-i-SIR
In Section 2.2, we employ the likelihood-free SIR technique iteratively to construct an efficient global kernel for ABC-MCMC. The proposal distribution plays a crucial role in importance sampling, influencing the efficiency of the sampling process. In essence, the proposal determines where samples are drawn from, and its similarity to the target distribution significantly impacts the quality of estimates. An effective proposal reduces variability and yields more reliable results. The calculation of weights depends on how well the proposal aligns with the target, and optimal performance is achieved when the proposal closely approximates the target distribution. In this section, we propose an adaptive procedure that iteratively enhances the importance proposal distribution within the ABC-i-SIR algorithm, aiming to progressively align it more closely with the target posterior distribution.
In this work, we leverage the idea of sequential importance sampling combined with a flexible density estimation framework based on weighted samples to progressively improve the proposal distribution. Specifically, we introduce a decreasing sequence of tolerance thresholds
$\varepsilon^{*}_{1}\geq\varepsilon^{*}_{2}\geq\cdots\geq\varepsilon^{*}_{T},$
and define a series of intermediate posterior distributions
$\pi_{\varepsilon^{*}_{1}}({\boldsymbol{\theta}}\mid\boldsymbol{y}),\;\pi_{\varepsilon^{*}_{2}}({\boldsymbol{\theta}}\mid\boldsymbol{y}),\;\ldots,\;\pi_{\varepsilon^{*}_{T}}({\boldsymbol{\theta}}\mid\boldsymbol{y}).$
Rather than directly approximating the target ABC posterior, we gradually update the proposal distribution to approach the target posterior through these intermediate distributions. The discrepancy between the proposal and the approximate target at each stage is reduced. This leads to a larger effective sample size of weighted samples and consequently leads to improved approximation quality. This series of intermediate target distribution is often used in ABC sequential Monte Carlo methods (del2012adaptive).
Concretely, the $j$-th update of the importance proposal distribution is constructed using the weighted sample set $\left\{\left({\boldsymbol{\theta}}_{(i)}^{*},w_{\varepsilon_{j}^{*},i}^{*}\right)\right\}_{i=1}^{n_{t}}$, where the normalized importance weights are given by
$w_{\varepsilon_{j}^{*},i}^{*}\propto\frac{\pi({\boldsymbol{\theta}}_{(i)}^{*})\,K_{\varepsilon_{j}^{*}}(\boldsymbol{x}_{(i)}^{*},\boldsymbol{y})}{q_{\varepsilon_{j-1}^{*}}^{*}({\boldsymbol{\theta}}_{(i)}^{*})}.$ (7)
Here $\{({\boldsymbol{\theta}}_{(i)}^{*},\boldsymbol{x}_{(i)}^{*})\}_{i=1}^{n_{t}}$ denotes a set of $n_{t}$ training data pairs sampled from the ABC-i-SIR algorithm, where the proposal distribution at the previous iterations is $q_{\varepsilon_{j-1}^{*}}$. The numerator, $\pi({\boldsymbol{\theta}}_{(i)}^{*})\,K_{\varepsilon_{j}^{*}}(\boldsymbol{x}_{(i)}^{*},\boldsymbol{y})$, represents the unnormalized intermediate target distribution $\pi_{\varepsilon_{j}^{*}}(({\boldsymbol{\theta}},\boldsymbol{x})\mid\boldsymbol{y})$.
To avoid increasing the simulation cost while ensuring a sufficient number of samples, we run a single Markov chain and update the importance proposal distribution once every $S$ ABC-i-SIR iterations. Specifically, $n_{t}=S\times N_{b}$ weighted samples are collected for each update, where $N_{b}$ denotes the batch size of each ABC-i-SIR iteration.
Kernel density estimation (KDE) (terrell1992variable) serves as a natural and computationally efficient choice for this density estimation step:
$q_{\varepsilon_{j}^{*}}^{*}({\boldsymbol{\theta}})=\sum_{i=1}^{n_{t}}w_{\varepsilon_{j}^{*},i}^{*}\,\tilde{K}_{h}({\boldsymbol{\theta}},{\boldsymbol{\theta}}_{(i)}^{*}),$
where $\tilde{K}_{h}$ represents a smooth kernel with bandwidth $h$. Nonetheless, KDE is not the sole viable estimator; alternative approaches such as Gaussian mixture models (mclachlan2014number), adaptive kernel methods (van2003adaptive), or generative models grounded in machine learning frameworks (prop-goodfellow2014generative; prop-kingma2013auto; papamakarios2021normalizing) can be employed to capture complex posterior geometries with potentially greater flexibility.
Moreover, to ensure the safety and robustness of the proposal distribution, an auxiliary tolerance threshold $\varepsilon^{*}_{T}$ can be set larger than the final target threshold $\varepsilon$. To guarantee the proposal distribution steadily and reliably approaches the target posterior, we adopt an adaptive scheme that gradually reduces the auxiliary tolerance while maintaining a sufficient number of nonzero-weighted training samples. Specifically, the sequence $\{\varepsilon^{*}_{j}\}$ is obtained by solving
$PA\left(\left\{\boldsymbol{x}_{(i)}^{*}\right\}_{i=1:n_{t}},\varepsilon^{*}_{j}\right)=\alpha\,PA\left(\left\{\boldsymbol{x}_{(i)}^{*}\right\}_{i=1:n_{t}},\varepsilon^{*}_{j-1}\right),$ (8)
where
$PA\left(\left\{\boldsymbol{x}_{(i)}^{*}\right\}_{i=1:n_{t}},\varepsilon^{*}_{j}\right)=\sum_{i=1}^{n_{t}}\mathbb{I}\left(\mathfrak{D}(\boldsymbol{x}_{(i)}^{*},\boldsymbol{y})<\varepsilon^{*}_{j}\right)$
denotes the number of alive particles under tolerance threshold $\varepsilon^{*}_{j}$. Here, $\mathbb{I}(\cdot)$ is the indicator function, $\mathfrak{D}(\boldsymbol{x},\boldsymbol{y})$ measures the discrepancy between the simulated data $\boldsymbol{x}$ and the observed data $\boldsymbol{y}$, and $0<\alpha<1$ controls the reduction rate of the proportion of alive samples. In general, the initial proposal distribution is set to the prior distribution, i.e., $q_{\varepsilon_{0}}=\pi$, with the initial tolerance threshold $\varepsilon_{0}=\infty$. In practice, it is recommended to set the parameter $\alpha$ within the range of 0.6 to 0.9. This range ensures a gradual reduction of the auxiliary threshold while maintaining a sufficiently large number of active particles to effectively update the proposal distribution. Algorithm 3 presents the detailed procedure for one iteration of the adaptive ABC-i-SIR (ABC-Ai-SIR) update.
Algorithm 3 One iteration of the ABC-Ai-SIR
[…]
Input: Previous state $({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$, simulator $P_{\boldsymbol{\theta}}$, kernel $K_{\varepsilon}$, prior $\pi({\boldsymbol{\theta}})$, number of steps to collect data $S$, auxiliary threshold $\varepsilon^{*}_{j}$, importance proposal density $q_{\varepsilon^{*}_{j}}$, target auxiliary threshold $\varepsilon^{*}_{T}$, auxiliary threshold reduction factor $\alpha$, currently collected training data $\mathcal{D}$.
Output: New state $({\boldsymbol{\theta}}_{t+1},\boldsymbol{x}_{{t+1}})$.
1: Set ${\boldsymbol{\theta}}_{(0)}^{*}={\boldsymbol{\theta}}_{t}$, and $\tilde{w}_{0}=\pi({\boldsymbol{\theta}}_{t})K_{\varepsilon}(\boldsymbol{x}_{t},\boldsymbol{y})/q_{j}({\boldsymbol{\theta}}_{t})$.
2:for $i=1:N_{b}$ do
3:  Draw ${\boldsymbol{\theta}}_{(1:N_{b})}^{*}$ from $p_{t}(\cdot)$ and generate simulate data $\boldsymbol{x}_{(i)}^{*}$ from $P_{{\boldsymbol{\theta}}_{(i)}^{*}}$.
4:  Compute weight $\tilde{w}_{i}=\pi({\boldsymbol{\theta}}_{(i)}^{*})K_{\varepsilon}(\boldsymbol{x}_{(i)}^{*},\boldsymbol{y})/q_{j}({\boldsymbol{\theta}}_{{(i)}}^{*})$.
5:end for
6: Sample ${\boldsymbol{\theta}}_{t+1}$ from $\left\{{\boldsymbol{\theta}}_{(i)}^{*}\right\}_{i=0:N_{b}}$ with weight $\left\{w_{i}\right\}_{i=0:N_{b}}$, where $w_{i}=\frac{\tilde{w}_{i}}{\sum_{i=1}^{N_{b}}\tilde{w}_{i}}$ for $i=1,\cdots,N_{b}$.
7: Collect $\left\{({\boldsymbol{\theta}}_{(i)}^{*},\boldsymbol{x}_{(i)}^{*})\right\}_{i=1:N_{b}}$ into the data set $\mathcal{D}$.
8:if $|D|=SN_{b}$ then
9:  $j=j+1$.
10:  If $\varepsilon^{*}_{j-1}>\varepsilon^{*}_{T}$, obtain $\varepsilon^{*}_{j}$ by solving Eq. (8), then set $\varepsilon^{*}_{j}=\max\{\varepsilon^{*}_{T},\varepsilon^{*}_{j}\}$.
11:  Obtain $q_{\varepsilon^{*}_{j}}$ by update $q_{\varepsilon^{*}_{j-1}}$ by using weighted data $\left\{\left({\boldsymbol{\theta}}_{(i)}^{*},w_{\varepsilon^{*}_{j},i}^{*}\right)\right\}_{i=1:SN_{b}}$ in the data set $\mathcal{D}$, where weight $w_{\varepsilon^{*}_{j},i}^{*}$ is defined in Eq. (7)
12:  Initialize $\mathcal{D}$ to an empty set.
13:end if
3.2 Sequential optimization of hyper-parameters